# Sparse VideoGen: Accelerating Video Diffusion Transformers with Spatial-Temporal Sparsity

### v1

@@ Abstract
Diffusion Transformers (DiTs) dominate video generation but their high computational cost severely limits real-world applicability, usually requiring tens of minutes to generate a few seconds of video even on high-performance GPUs. This inefficiency primarily arises from the quadratic computational complexity of 3D full attention with respect to the context length. In this paper, we propose a training-free framework termed Sparse VideoGen (SVG) that leverages the inherent sparsity in 3D full attention to boost inference efficiency. We reveal that the attention heads can be dynamically classified into two groups depending on distinct sparse patterns: (1) Spatial Head, where only spatially-related tokens within each frame dominate the attention output, and (2) Temporal Head, where only temporally-related tokens across different frames dominate. Based on this insight, SVG proposes an online profiling strategy to capture the dynamic sparse patterns and predicts the type of attention head. Combined with a novel hardware-efficient tensor layout transformation and customized kernel implementations, SVG achieves up to 2.28×2.28$\times$2.28 × and 2.33×2.$3$ end-to-end speedup on CogVideoX-v1.5 and HunyuanVideo, respectively, while preserving generation quality.
@@ 1 Introduction
Diffusion Transformers (DiTs) (Peebles & Xie, 2023) have recently emerged as a transformative paradigm for generative tasks, achieving state-of-the-art results in image generation. This success has been naturally carried over to video generation, with models adapting from a spatial 2D attention to a spatiotemporal 3D full attention (Arnab et al., 2021; Yang et al., 2024c; Kong et al., 2024), resulting in high-fidelity and temporally consistent outputs. Close-sourced models such as Sora and Kling, and open-sourced models including CogVideo (Hong et al., 2023) and HunyuanVideo (Kong et al., 2024), have showcased impressive capabilities in applications ranging from animation (Guo et al., 2024) to physical world simulation (Liu et al., 2024b).
@@
Fortunately, attention in transformers is well-known for its sparsity, offering a great opportunity to reduce redundant computation. For example, in large language models (LLMs), a small portion of the tokens can dominate the attention output (Zhang et al., 2023b; Xiao et al., 2024b; Tang et al., 2024). Therefore, the computation can be dramatically reduced by only computing the attention among such important tokens, while still maintaining generation accuracy. However, existing methods cannot be directly applied to DiTs (as shown in Table 1), as video data has fundamentally different sparsity patterns from text data.
@@
  • We conduct in-depth analyses of video DiTs’ sparse patterns, revealing two inherent sparse attention patterns (temporal head and spatial head) for efficient video generation.
@@ 2.1 Efficient diffusion models
Diffusion model compression. Weight compression through quantization is a common tactic (Li et al., 2023; Zhao et al., 2024a; Li* et al., 2025), pushing attention modules to INT8 (Zhang et al., 2025) or even INT4/FP8 (Zhang et al., 2024). Other work proposes efficient architectures (Xie et al., 2024; Cai et al., 2024; Chen et al., 2025) or high-compression autoencoders (Chen et al., 2024a) to improve performance. Our Sparse VideoGen is orthogonal to these techniques and can incorporate them for additional gains.
@@ 2.2 Efficient attention methods
Sparse attention in LLMs. Recent research on sparse attention in language models reveals diverse patterns to reduce computational overhead. StreamingLLM (Xiao et al., 2023) and LM-Infinite (Han et al., 2023) observe that attention scores often concentrate on the first few or local tokens, highlighting temporal locality. H2O (Zhang et al., 2023a), Scissorhands (Liu et al., 2024d) and DoubleSparsity (Yang et al., 2024b) identify a small set of “heavy hitter” tokens dominating overall attention scores. TidalDecode (Yang et al., 2024a) shows that attention patterns across layers are highly correlated, while DuoAttention (Xiao et al., 2024a) and MInference (Jiang et al., 2024) demonstrate distinct sparse patterns across different attention heads. However, these methods focus on token-level sparsity and do not leverage the inherent redundancy of video data.
Linear and low-bit attention. Another direction involves linear attention (Cai et al., 2023; Xie et al., 2024; Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020), which lowers complexity from quadratic to linear, and low-bit attention (Zhang et al., 2025, 2024), which operates in reduced precision to accelerate attention module. Sparse VideoGen is orthogonal to both approaches: it can be combined with techniques like FP8 attention while still benefiting from the video-specific spatial and temporal sparsity in video diffusion models.
@@ 4.1 Online profiling strategy for sparsity identification
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ VBench-1 $\uparrow$ VBench-2 $\uparrow$ FLOPS $\downarrow$ Latency $\downarrow$ Speedup $\uparrow$
@@ 4.3 Other optimizations
Quantization. We further incorporate FP8 quantization into sparse attention (Zhang et al., 2025, 2024; Zhao et al., 2024c), which further boosts up to 1.3×1.3$\times$1.3 × throughput with minimal accuracy drop as shown in Table 1. We also customize an attention kernel that supports both FP8 quantization and block sparse computation.
@@ 5.1 Setup
Metrics. We assess the quality of the generated videos using the following metrics. We use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Structural Similarity Index Measure (SSIM) to evaluate the generated video’s similarity, and use VBench Score (Huang et al., 2023) to evaluate the video quality, following common practices in community (Horé & Ziou, 2010; Zhao et al., 2024b; Li* et al., 2025; Li et al., 2024). Specifically, we report the imaging quality and subject consistency metrics, represented by VBench-1 and VBench-2 in our table.
@@ A.1 Efficient Diffusion Models
Diffusion Models function primarily as denoising models that are trained to estimate the gradient of the data distribution (Song & Ermon, 2019). Although these models are capable of generating samples with high quality and diversity, they are known as inefficient. To enhance the efficiency of diffusion models, researchers often focus on three primary approaches: (1) decreasing the number of denoising steps, (2) reducing the model size, and (3) optimizing system implementation for greater efficiency.
@@
A common approach to enhancing the efficiency of diffusion models involves compressing their weights through quantization. Q-Diffusion (Li et al., 2023) introduced a W8A8 strategy, implementing quantization in these models. Building on this foundation, ViDiT-Q (Zhao et al., 2024a) proposed a timestep-aware dynamic quantization method that effectively reduces the bit-width to W4A8. Furthermore, SVDQuant (Li* et al., 2025) introduced a cost-effective branch designed to address outlier problems in both activations and weights, thus positioning W4A4 as a feasible solution for diffusion models. SageAttention (Zhang et al., 2025) advanced the field by quantizing the attention module to INT8 precision via a smoothing technique. SageAttention V2 (Zhang et al., 2024) extended these efforts by pushing the precision boundaries to INT4 and FP8. Another common approach is to design efficient diffusion model architectures (Xie et al., 2024; Cai et al., 2024; Chen et al., 2025) and high-compression autoencoders (Chen et al., 2024a) to boost efficiency. Our Sparse VideoGen is orthogonal to these techniques and can utilize them as supplementary methods to enhance efficiency.
@@ A.2 Efficient Attention
Recent studies on sparse attention in language models have identified patterns that reduce computational costs by targeting specific token subsets. StreamingLLM (Xiao et al., 2023) and LM-Infinite (Han et al., 2023) reveal concentration on initial and local tokens, highlighting temporal locality in decoding. H2O (Zhang et al., 2023a) and Scissorhands (Liu et al., 2024d) note attention focuses mainly on a few dominant tokens. TidalDecode (Yang et al., 2024a) shows cross-layer attention pattern correlations, aiding in attention sparsity. DuoAttention (Xiao et al., 2024a) and MInference (Jiang et al., 2024) find distinct sparse patterns among attention heads, with varying focus on key tokens and context. Despite their success in LLMs, these mechanisms are constrained to token-level sparsity and miss the redundancy unique to video data.
@@
Significant advancements have been achieved in enhancing attention efficiency, notably through linear attention (Cai et al., 2023; Xie et al., 2024) and low-bit attention techniques (Zhang et al., 2025, 2024). Linear attention models, including Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), MetaFormer (Yu et al., 2022), and LinearAttention (Katharopoulos et al., 2020), reduce the quadratic complexity of traditional attention to linear. Low-bit attention approaches decrease computational demands by utilizing lower precision, with SageAttention (Zhang et al., 2025) employing INT8 precision to enhance efficiency without notable performance loss.

### v2

@@ Abstract
[…]
@@ 2 Introduction
Diffusion Transformers (DiTs) (Peebles & Xie, 2023) have recently emerged as a transformative paradigm for generative tasks, achieving state-of-the-art results in image generation. This success has been naturally carried over to video generation, with models adapting from a spatial 2D attention to a spatiotemporal 3D full attention (Arnab et al., 2021; Yang et al., 2024c; Kong et al., 2024), resulting in high-fidelity and temporally consistent outputs. Close-sourced models such as Sora and Kling, and open-sourced models including Wan 2.1 (Wang et al., 2025), CogVideo (Hong et al., 2023), and HunyuanVideo (Kong et al., 2024), have showcased impressive capabilities in applications ranging from animation (Guo et al., 2024; Feng et al., 2024) to physical world simulation (Liu et al., 2024b).
@@
Fortunately, attention in transformers is well-known for its sparsity, offering a great opportunity to reduce redundant computation. For example, in large language models (LLMs), a small portion of the tokens can dominate the attention output (Zhang et al., 2023c; Xiao et al., 2024b; Tang et al., 2024). Therefore, the computation can be dramatically reduced by only computing the attention among such important tokens, while still maintaining generation accuracy. However, existing methods cannot be directly applied to DiTs (as shown in Table 1), as video data has fundamentally different sparsity patterns from text data.
@@
  • We conduct in-depth analysis of video DiTs’ sparse patterns, revealing two inherent sparse attention patterns (temporal head and spatial head) for efficient video generation.
@@ 3.1 Efficient diffusion models
Diffusion model compression. Weight compression through quantization is a common tactic (Li et al., 2023; Zhao et al., 2024a; Li* et al., 2025), pushing attention modules to INT8 (Zhang et al., 2025a) or even INT4/FP8 (Zhang et al., 2024). Other work proposes efficient architectures (Xie et al., 2024; Cai et al., 2024; Chen et al., 2025) or high-compression autoencoders (Chen et al., 2024a) to improve performance. Our Sparse VideoGen is orthogonal to these techniques and can incorporate them for additional gains.
@@ 3.2 Efficient attention methods
Sparse attention in LLMs. Recent research on sparse attention in language models reveals diverse patterns to reduce computational overhead. StreamingLLM (Xiao et al., 2023) and LM-Infinite (Han et al., 2023) observe that attention scores often concentrate on the first few or local tokens, highlighting temporal locality. H2O (Zhang et al., 2023b), Scissorhands (Liu et al., 2024d) and DoubleSparsity (Yang et al., 2024b) identify a small set of “heavy hitter” tokens dominating overall attention scores. TidalDecode (Yang et al., 2024a) shows that attention patterns across layers are highly correlated, while DuoAttention (Xiao et al., 2024a) and MInference (Jiang et al., 2024) demonstrate distinct sparse patterns across different attention heads. However, these methods focus on token-level sparsity and do not leverage the inherent redundancy of video data.
Linear and low-bit attention. Another direction involves linear attention (Cai et al., 2023; Xie et al., 2024; Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020), which lowers complexity from quadratic to linear, and low-bit attention (Zhang et al., 2025a, 2024), which operates in reduced precision to accelerate attention module. Sparse VideoGen is orthogonal to both approaches: it can be combined with techniques like FP8 attention while still benefiting from the video-specific spatial and temporal sparsity in video diffusion models.
@@ 5.1 Online profiling strategy for sparsity identification
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ImageQual $\uparrow$ SubConsist $\uparrow$ FLOPS $\downarrow$ Latency $\downarrow$ Speedup $\uparrow$
@@ 5.3 Other optimizations
Quantization. We further incorporate FP8 quantization into sparse attention (Zhang et al., 2025a, 2024; Zhao et al., 2024c), which further boosts up to 1.3×1.3$\times$1.3 × throughput with minimal accuracy drop as shown in Table 1. We also customize an attention kernel that supports both FP8 quantization and block sparse computation.
@@ 6.1 Setup
Metrics. We assess the quality of the generated videos using the following metrics. We use Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Structural Similarity Index Measure (SSIM) to evaluate the generated video’s similarity, and use VBench Score (Huang et al., 2023) to evaluate the video quality, following common practices in community (Horé & Ziou, 2010; Zhao et al., 2024b; Li* et al., 2025; Li et al., 2024). Specifically, we report the imaging quality and subject consistency metrics, represented by ImageQaul and SubConsist in our table.
@@
Visualizations. We present a comparison of the videos generated by Dense Attention and Sparse VideoGen in Appendix B. Additionally, real video samples are available on Google Drive and can be accessed here.
@@ A.1 Efficient Diffusion Models
Diffusion Models function primarily as denoising models that are trained to estimate the gradient of the data distribution (Song & Ermon, 2019; Zhang et al., 2023a). Although these models are capable of generating samples with high quality and diversity, they are known as inefficient. To enhance the efficiency of diffusion models, researchers often focus on three primary approaches: (1) decreasing the number of denoising steps, (2) reducing the model size, and (3) optimizing system implementation for greater efficiency.
@@
A common approach to enhancing the efficiency of diffusion models involves compressing their weights through quantization. Q-Diffusion (Li et al., 2023) introduced a W8A8 strategy, implementing quantization in these models. Building on this foundation, ViDiT-Q (Zhao et al., 2024a) proposed a timestep-aware dynamic quantization method that effectively reduces the bit-width to W4A8. Furthermore, SVDQuant (Li* et al., 2025) introduced a cost-effective branch designed to address outlier problems in both activations and weights, thus positioning W4A4 as a feasible solution for diffusion models. SageAttention (Zhang et al., 2025a) advanced the field by quantizing the attention module to INT8 precision via a smoothing technique. SageAttention V2 (Zhang et al., 2024) extended these efforts by pushing the precision boundaries to INT4 and FP8. Another common approach is to design efficient diffusion model architectures (Xie et al., 2024; Cai et al., 2024; Chen et al., 2025) and high-compression autoencoders (Chen et al., 2024a) to boost efficiency. Our Sparse VideoGen is orthogonal to these techniques and can utilize them as supplementary methods to enhance efficiency.
@@ A.2 Efficient Attention
Recent studies on sparse attention in language models have identified patterns that reduce computational costs by targeting specific token subsets. StreamingLLM (Xiao et al., 2023) and LM-Infinite (Han et al., 2023) reveal concentration on initial and local tokens, highlighting temporal locality in decoding. H2O (Zhang et al., 2023b) and Scissorhands (Liu et al., 2024d) note attention focuses mainly on a few dominant tokens. TidalDecode (Yang et al., 2024a) shows cross-layer attention pattern correlations, aiding in attention sparsity. DuoAttention (Xiao et al., 2024a) and MInference (Jiang et al., 2024) find distinct sparse patterns among attention heads, with varying focus on key tokens and context. MMInference  (Li et al., 2025) speedup the vision language models through modality-aware permutation. SpargeAttention (Zhang et al., 2025b) and XAttention (Xu et al., 2025) propose general sparsity identification algorithms that can be applied to all forms of models. Despite their success in LLMs or VLMs, these mechanisms are constrained to token-level sparsity and miss the redundancy unique to video data.
@@
Significant advancements have been achieved in enhancing attention efficiency, notably through linear attention (Cai et al., 2023; Xie et al., 2024) and low-bit attention techniques (Zhang et al., 2025a, 2024). Linear attention models, including Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), MetaFormer (Yu et al., 2022), and LinearAttention (Katharopoulos et al., 2020), reduce the quadratic complexity of traditional attention to linear. Low-bit attention approaches decrease computational demands by utilizing lower precision, with SageAttention (Zhang et al., 2025a) employing INT8 precision to enhance efficiency without notable performance loss.
@@
Appendix B Visualization of the generated videos
We provide visualization comparison between Dense Attention and Sparse VideoGen on HunyuanVideo and Wan 2.1. We conduct both Text-to-Video generation and Image-to-Video generation under 720p resolution. Results demonstrates that Sparse VideoGen can preserve high pixel-level fidelity, achieving similar generation quality compared with the dense attention.
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Figure 9: Comparion of Dense Attention and Sparse VideoGen on HunyuanVideo Text-to-Video generation.
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Figure 10: Comparion of Dense Attention and Sparse VideoGen on Wan 2.1 Text-to-Video generation.
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Dense Attention
Sparse VideoGen
Figure 11: Comparison of Dense Attention and Sparse VideoGen on Wan 2.1 Image-to-Video generation.